OpenStudy (anonymous):

Can someone check my work plz? 2/3n=8 n=____ I get 9 Is that right? IS the correct answer n=12???

OpenStudy (anonymous):

how did u get that ?

OpenStudy (anonymous):

tbh I took 2/3 and multiplied to get a whole # 2/383=6/3 which is the whole #2

OpenStudy (anonymous):

hold on ...think I goofed on it

OpenStudy (anonymous):

waiiit a min i am confused see the easiest method

OpenStudy (anonymous):

so frustrated.... now i can not duplicate the problem... ugh

OpenStudy (anonymous):

2 = 3n* 8 do it like this

OpenStudy (anonymous):

2 = 24n

OpenStudy (anonymous):

so n 2/24

OpenStudy (anonymous):

n = 1 /12

OpenStudy (anonymous):

i understand the problem as 2/3 multiplied by N gives you the answer of 8 therefor 2/3 goes into 8 ____ times the ____ being what N= right?

OpenStudy (anonymous):

nooo dont confuse it take it simpleeee

OpenStudy (anonymous):

2/3n = 8\[\frac{ 2 }{ 3n } =8\]

OpenStudy (anonymous):

therefore 2 = 3n * 8

OpenStudy (anonymous):

2 = 24n

OpenStudy (anonymous):

24n = 2

OpenStudy (anonymous):

n = 2/24

OpenStudy (anonymous):

n = 1 /12

OpenStudy (anonymous):

thats it

OpenStudy (anonymous):

i just got n=12 using the http://www.tiger-algebra.com/drill/2/3n=8/

OpenStudy (anonymous):

Is it 2/3n or (2/3)*n?

OpenStudy (anonymous):

no its -8 and in ur equation it is =8

OpenStudy (anonymous):

it says solve. 2over 3n=8 n=___ so 2/3n=8 n=__

OpenStudy (anonymous):

So then Smitha is right

OpenStudy (anonymous):

idk.... I went to http://www.tiger-algebra.com/ and used their system and it says only one answer can be found and that is n=12. looking at how it is worked out it appears to be correct but now i am second guessing things. I only have one try with this test so I can not afford to fail it :/

OpenStudy (anonymous):

If I could suggest something real quick, your equation is 2/3n = 8. If you have a calculator on you you could find that 2/3 is .6 (repeated). So .6n(repeated) = 8. Then you have to divide the .6n (repeated) by .6 to get "n" by itself. And the 8 by the .6 as well. The answer I got was 12. Sorry if this didn't help.

OpenStudy (anonymous):

Thank you. It helps and that is the same principle steps I was doing

OpenStudy (anonymous):

Ok, and no problem :D This used to be my favorite type of math.

OpenStudy (anonymous):

can you maybe help me with a few more , if you have time?

OpenStudy (anonymous):

Sure. Just give me a minute or two with my responses :P

OpenStudy (anonymous):

after an hour and a half i am only on #5...lol...

OpenStudy (anonymous):

There is a difference between 2/3n and and (2/3)*n

OpenStudy (anonymous):